A cyclist and his bike have a combine mass of 75kg. Find the maximum speed that the cyclist can attain on a level ground when working at a constant rate of 210W. Against a resistance of 21N, with the resistance and rate of working kept constant the cyclist attain an inclination 3.8°. Find the maximum speed of the cyclist up the slope.


1
Expert's answer
2021-01-27T15:28:28-0500

1)

"P=Fv=mgv," "\u21d2" "v=\frac{P}{mg},"

"v=\frac{210}{75\cdot 9.8}\approx 0.286 ~\frac ms."

2)

"P=(\frac{F}{cos\alpha}+F')v," "\u21d2" "v=\frac{P}{\frac{mg}{cos\alpha}+F'},"

"v=\frac{210}{\frac{75\cdot 9.8}{cos 3.8°}+21}\approx0.277 ~\frac ms."